Novak Company purchased a new machine on October 1, 2020, at a cost of $110,100. The company estimated that the machine will have a salvage value of $10,100. The machine is expected to be used for 10,000 working hours during its 4-year life.

(a)

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Compute the depreciation expense under straight-line method for 2020. (Round answer to 0 decimal places, e.g. 2,125.)

2020

Depreciation expense

$enter a dollar amount rounded to 0 decimal places
Answer #1

Solution

2020
Depreciation expense $ 6,250

Working

Straight line Method
A Cost $ 110,100
B Residual Value $ 10,100
C=A - B Depreciable base $ 100,000
D Life [in years left ]                                  4
E=C/D Annual SLM depreciation $ 25,000.00

Depreciation for a year is 25000 but in 2020 depreciation will be recorded for only 3 months.......25000/12 x 3=$6250